What are the graph line coordinates for x + y = 7 and 3x + 4y = 24
Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

y=-x+7
(0,7)
(1,6)
(2,5)
-------------
3x+4y=24
4y=-3x+24
y=-3/4x+6
(0,6)
(4,3)
(8,0)

Answer by richwmiller(17219)